# Building temporally coherent 3D maps with VGGT for memory-efficient Semantic SLAM

### v1

@@ Abstract
We present a fast, spatio-temporal scene understanding framework based on Vision Gated Generative Transformers (VGGT). The proposed pipeline is designed to enable efficient, close to real-time performance, supporting applications including assistive navigation. To achieve continuous updates of the 3D scene representation, we process the image flow with a sliding window, aligning submaps, thereby overcoming VGGT’s high memory demands. We exploit the VGGT tracking head to aggregate 2D semantic instance masks into 3D objects. To allow for temporal consistency and richer contextual reasoning the system stores timestamps and instance-level identities, thereby enabling the detection of changes in the environment. We evaluate the approach on well-known benchmarks and custom datasets specifically designed for assistive navigation scenarios. The results demonstrate the applicability of the framework to real-world scenarios.
@@ 2 Related work
VGGT Long [2] extends VGGT’s memory-limited, locally consistent point-map generation to kilometer-scale scenes. It partitions the sequence into overlapping chunks (6̃0-75 frames), aligns each chunk to the previous via Iteratively Reweighted Least Squares (IRLS) optimization, and mitigates drift through loop detection and Sim(3) alignment of overlapping pointclouds, followed by a global optimization over the chunk transforms. Because it requires multiple passes and access to the full sequence—including a final global adjustment over all frames—the method is not suited to streaming input; moreover, dynamic objects are typically down-weighted and effectively treated as outliers due to low confidence.

### v2

@@ Abstract
We present a fast, spatio-temporal scene understanding framework based on Visual Geometry Grounded Transformer (VGGT). The proposed pipeline is designed to enable efficient, close to real-time performance, supporting applications including assistive navigation. To achieve continuous updates of the 3D scene representation, we process the image flow with a sliding window, aligning submaps, thereby overcoming VGGT’s high memory demands. We exploit the VGGT tracking head to aggregate 2D semantic instance masks into 3D objects. To allow for temporal consistency and richer contextual reasoning the system stores timestamps and instance-level identities, thereby enabling the detection of changes in the environment. We evaluate the approach on well-known benchmarks and custom datasets specifically designed for assistive navigation scenarios. The results demonstrate the applicability of the framework to real-world scenarios.
@@ 2 Related work
VGGT Long [2] extends VGGT’s memory-limited, locally consistent point-map generation to kilometer-scale scenes. It partitions the sequence into overlapping chunks ($\sim$60-75 frames), aligns each chunk to the previous via Iteratively Reweighted Least Squares (IRLS) optimization, and mitigates drift through loop detection and Sim(3) alignment of overlapping pointclouds, followed by a global optimization over the chunk transforms. Because it requires multiple passes and access to the full sequence—including a final global adjustment over all frames—the method is not suited to streaming input; moreover, dynamic objects are typically down-weighted and effectively treated as outliers due to low confidence.
@@ 4.5 Conclusion
5 Acknowledgments
The work was supported by the TKP2021-NVA-27 grant, funded by the Ministry of Innovation and Technology with support from the National Research Development and Innovation Fund under the TKP2021 program. Gergely Dinya was supported by the Research Fellowship EKÖP-25-2-I-ELTE-987 of the Ministry of Culture and Innovation of Hungary, funded by the National Fund for Research, Development and Innovation (NRDI).